OpenStudy (anonymous):

can you help with this one? Solve the equation. (If there is no solution, enter NO SOLUTION.) |x − 7| = |3x + 8|

OpenStudy (anonymous):

you have to solve two separate equations, namely \[x-7=3x+8\] and \[x-7=-(3x+8)\]

OpenStudy (anonymous):

in other words just remove absolute value signs for one equation, then for the next take the negative of one of the sides (doesn't matter which)

OpenStudy (anonymous):

first one is easy enough \[x-7=3x+8\][ \[-7-8=3x-x\] \[-15=2x\] \[x=-\frac{15}{2}\]

OpenStudy (anonymous):

second one first distribute the minus sign to get \[x-7=-3x-8\] then proceed as usual \[x+3x=-8+7\] \[4x=-1\] \[x=-\frac{1}{4}\]

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

got it so. -1/4 is obviously the larger value

OpenStudy (anonymous):

yes for sure